OpenStudy (didee):

I cannot get the format right for the vectors, sorry: Suppose "underscore" u = (u1, u2, u3) Is this true? "underscore" u "dot" u = sqrt(u1^2 + u2^2 + u3^2 Please explain in detail, thanks

OpenStudy (amistre64):

try it out.

OpenStudy (amistre64):

<a,b,c>.<x,y,z> = ax + by + cz right?

OpenStudy (didee):

ok.....

OpenStudy (amistre64):

whats an underscore u?

OpenStudy (didee):

like i said, can't get the format right sorry. I'm referring to a vector. I have a solution but i dont understand it, will try and post it now

OpenStudy (amistre64):

since we are taking vectors, itll be assumed that u is a vector

OpenStudy (amistre64):

perhaps the notation: u = <u1,u2,u3> might make life simpler

OpenStudy (amistre64):

i have a meeting starting soon, so ill check in on this in about an hour to see if anyones gotten back to you

OpenStudy (didee):

I've attached, thanks.

OpenStudy (didee):

I have more questions on vectors and de moivre's theorem. Will mail it to you if you don' t mind. Thanks

OpenStudy (amistre64):

i notice that the underscore is your materials way of expressing a vector then. Alot of texts boldface it or put an arrow above it, but all that is for is trying to come up with a suitable means of getting a point across. the dot product is an operation on vectors that is related to multiplication, except for in R^2 instead of R^1 a vector of R^1 would just look like a number, an ordered pair with a single element, u = 3 would be a vector from the origin to the point x=3. v = -2 would be a vector from the origin to the point x=-2 u.v is defined the same way as other dimensional ordered pairs, but since there is no other element in the setup, its just multiplication of like elements, u = 3 v = -2 -------- u.v = -6 When we lift this into higher dimensions, we have the same concepts, but after we multiply like elements, we add up all the results. say: u = < 3,5> v = <-2,2> ------------ u.v = -6+10 = 4

OpenStudy (didee):

Hi, thanks so much for your explanation. It is much appreciated.